# ON EQUICONTINUITY OF MAPPINGS WITH VARIABLE DEFINITION DOMAINS

### v1

@@ 3 On image of the continuum under homeomorphisms
Indeed, let $\gamma:[0,1]\rightarrow\overline{{\mathbb{R}}^{n}}$ be a path in $\Gamma(f_{m}(A_{m}),E_{m},\overline{{\mathbb{R}}^{n}}),$ i.e., $\gamma(0)\in f_{m}(A_{m}),$ $\gamma(1)\in E_{m}$ and $\gamma(t)\in\overline{{\mathbb{R}}^{n}}$ for $t\in(0,1).$ Let $t_{m}=\sup\limits_{\gamma(t)\in f_{m}(D_{m})}t$ and let $\alpha_{m}(t)=\gamma|_{[0,t_{m})}.$ Let $\Gamma_{m}$ consists of all such paths $\alpha_{m},$ now $\Gamma(f_{m}(A_{m}),E_{m},\overline{{\mathbb{R}}^{n}})>\Gamma_{m}$ and by the minorization principle of the modulus (see [Fu, Theorem 1])
$M(\Gamma_{m})\geqslant M(\Gamma(f_{m}(A_{m}),E_{m},\overline{{\mathbb{R}}^{n}}))\,.$ (3.4)
Let $\Delta_{m}$ be a family of all total $f_{m}$ liftings of $\Gamma_{m}$ starting at $A_{m}$ (it exists by Proposition 3). Now
@@
Let $A_{m}\subset\bigcup\limits_{i=1}^{N}\overline{B(z_{mi},r_{mi})},$ $r_{mi}>0,$ where $z_{mi}\in A_{m},$ $1\leqslant i\leqslant N$ and $h(\overline{B(z_{mi},r_{mi})},\partial D_{m})\geqslant\delta.$ By the compactness of $\overline{{\mathbb{R}}^{n}},$ we may consider that $z_{mi}\rightarrow z_{0i}$ as $m\rightarrow\infty.$
Since $h(\overline{B(z_{mi},r_{mi})},\partial D_{m})\geqslant\delta,$ we have that $B_{h}(z_{mi},\delta/2)\subset D_{m}$ for every $m=1,2,\ldots$ and $1\leqslant i\leqslant N,$ where $B_{h}(w_{0},p):=\{x\in\overline{{\mathbb{R}}^{n}}:h(x,w_{0})<p\}.$ Indeed, otherwise $B_{h}(z_{mi},\delta/2)\cap D_{m}\neq\varnothing\neq B_{h}(z_{mi},\delta/2)\setminus D_{m}.$ Since $B_{h}(z_{mi},\delta/2)$ is connected, we have that $B_{h}(z_{mi},\delta/2)\cap\partial D_{m}\neq\varnothing.$ Now, $h(z_{mi},\partial D_{m})\leqslant h(\overline{B(z_{mi},r_{mi})},\partial D_{m})\leqslant\delta/2,$ which contradicts with the condition $h(\overline{B(z_{mi},r_{mi})},\partial D_{m})\geqslant\delta.$ Thus, $h(\overline{B(z_{mi},r_{mi})},\partial D_{m})\geqslant\delta,$ as required.
Since $z_{mi}\rightarrow z_{0i}$ as $m\rightarrow\infty,$ we have that $h(z_{mi},z_{0i})<\delta/4$ for sufficiently large $m\in{\mathbb{N}}.$ Now, $z_{0i}\in B_{h}(z_{mi},\delta/2)\subset D_{m}$ for above $m.$ In particular, it follows from the mentioned above that $z_{0i}\neq\infty.$ We may consider that, $z_{0i}\in D_{m}$ for all $m\in{\mathbb{N}}.$
In the similar way, we may consider that $r_{mi}\rightarrow r_{0i}$ as $m\rightarrow\infty$ for some $r_{0i}\in{\mathbb{R}}^{\,+}\cup\{\infty\}.$ Observe that $r_{0i}\neq\infty.$ Otherwise, there is $w_{m}\in B(z_{mi},r_{mi}),$ $w_{m}\in S(z_{mi},r_{mi}/2),$ where $z_{mi}\rightarrow z_{0i}\neq\infty$ as $m\rightarrow\infty,$ and $r_{mi}\rightarrow\infty$ as $m\rightarrow\infty.$ Now, by the triangle inequality,
$|w_{m}|\geqslant|w_{m}-z_{mi}|-|z_{mi}|\geqslant r_{mi}/2-c\rightarrow\infty\,$,
$m\rightarrow\infty,$ where $c>0$ is some constant. Thus, $w_{m}\in B(z_{mi},r_{mi})$ and $w_{m}\rightarrow\infty$ as $m\rightarrow\infty.$ Now, $h(\overline{B(z_{mi},r_{mi})},\infty)\rightarrow 0$ as $m\rightarrow\infty.$ The latter contradicts the condition $h(\overline{B(z_{mi},r_{mi})},\partial D_{m})\geqslant\delta.$ Indeed, by Proposition 3
$h(\overline{B(z_{mi},r_{mi})},\infty)\geqslant h(\overline{B(z_{mi},r_{mi})},\partial D_{m})\geqslant\delta\,.$
The latter proves that $r_{0i}\neq\infty.$
Let $x\in\overline{B(z_{mi},r_{mi})},$ $r_{mi}\geqslant 0.$ Now, by the triangle inequality
$|x-z_{0i}|\leqslant|x-z_{mi}|+|z_{mi}-z_{0i}|<r_{0i}+\delta_{mi}\,$, (3.6)
where $\delta_{mi}>0,$ $\delta_{mi}\rightarrow 0$ as $m\rightarrow\infty.$ The relation (3.6) implies that
$\overline{B(z_{mi},r_{mi})}\subset B(z_{0i},r_{0i}+\delta_{mi})\,.$ (3.7)
Observe that,
$\overline{B(z_{0i},r_{0i})}\subset D_{m}$ (3.8)
for sufficiently large $m\in{\mathbb{N}}.$ Indeed, let $\alpha_{mi}=h(\overline{B(z_{0i},r_{0i})},\partial D_{m})=h(a_{mi},b_{mi}),$ where $a_{mi}\in\overline{B(z_{0i},r_{0i})}$ and $b_{mi}\in\partial D_{m}.$ Let $h(\overline{B(z_{mi},r_{mi})},\partial D_{m})=h(c_{mi},d_{mi})\geqslant\delta>0,$ $c_{mi}\in\overline{B(z_{mi},r_{mi})},$ $d_{mi}\in\partial D_{m}.$ Now, by the triangle inequality
$\alpha_{mi}=h(a_{mi},b_{mi})\geqslant h(b_{mi},c_{mi})-h(c_{mi},a_{mi})\geqslant$
$\geqslant\delta-|c_{mi}-a_{mi}|\geqslant\delta-|c_{mi}-z_{mi}|-|z_{mi}-z_{0i}|-|z_{0i}-a_{mi}|$ (3.9)
$\geqslant\delta-r_{mi}-r_{0i}-|z_{mi}-z_{0i}|\geqslant\delta/2$
for sufficiently large $m\in{\mathbb{N}}.$ Now, (3.9) proves (3.8). Since $\delta_{mi}\rightarrow 0$ as $m\rightarrow\infty$ in (3.6), the relation (3.8) implies that we may find some $r^{i}_{2}>r^{i}_{1}>r_{0i}>0$ such that $r_{0i}+\delta_{mi}<r^{i}_{1}$ and simultaneously $\overline{B(z_{0i},r^{i}_{2})}\subset D_{m}$ for sufficiently large $m.$ Now, (3.7) implies that $A_{m}\subset\bigcup\limits_{i=1}^{N}B(z_{0i},r^{i}_{1}).$ We now show that
$\Delta_{m}>\bigcup\limits_{i=1}^{N}\Gamma(S(z_{0i},r^{i}_{1}),S(z_{0i},r^{i}_{2}),A(z_{0i},r^{i}_{1},r^{i}_{2}))\,.$ (3.10)
@@
Indeed, let $\gamma\in\Delta_{m},$ in other words, $\gamma:[0,t_{m})\rightarrow D_{m},$ $\gamma(0)\in A_{m}$ and $\gamma(t)\rightarrow\partial D_{m}$ as $t\rightarrow t_{m}-0.$ Now, there is $1\leqslant i\leqslant N$ such that $|\gamma|\cap B(z_{0i},r^{i}_{1})\neq\varnothing\neq|\gamma|\cap(D_{m}\setminus B(z_{0i},r^{i}_{1})).$ Therefore, by Proposition 3 there is $0<t^{\,*}_{1}<t_{m}$ such that $\gamma(t^{\,*}_{1})\in S(z_{0i},r^{i}_{1}).$ We may assume that $\gamma(t)\not\in B(z_{0i},r^{i}_{1})$ for $t>t^{\,*}_{1}.$ Put $\gamma_{1}:=\gamma|_{[t^{\,*}_{1},t_{m}]}.$ Similarly, $|\gamma_{1}|\cap B(z_{0i},r^{i}_{2})\neq\varnothing\neq|\gamma_{1}|\cap(D_{m}\setminus B(z_{0i},r^{i}_{2})).$ By Proposition 3 there is $t^{\,*}_{1}<t^{\,*}_{2}<t_{m}$ with $\gamma(t^{\,*}_{2})\in S(z_{0i},r^{i}_{2}).$ We may assume that $\gamma(t)\in B(z_{0i},r^{i}_{2})$ for $t<t^{\,*}_{2}.$ Put $\gamma_{2}:=\gamma|_{[t^{\,*}_{1},t^{\,*}_{2}]}.$ Then, the path $\gamma_{2}$ is a subpath of $\gamma,$ which belongs to the family $\Gamma(S(z_{0i},r^{i}_{1}),S(z_{0i},r^{i}_{2}),A(z_{0i},r^{i}_{1},r^{i}_{2})).$ Thus, the relation (3.10) is established.
@@
$\psi_{i}(t)=\left\{\begin{array}[]{rr}1/[tq_{z_{0i}}^{\frac{1}{n-1}}(t)]\ ,&\ t\in(r^{i}_{1},r^{i}_{2})\ ,\\ 0\ ,&\ t\notin(r^{i}_{1},r^{i}_{2})\ .\end{array}\right$.
Now,
$\int\limits_{A(z_{0i},r^{i}_{1},r^{i}_{2})}Q(x)\cdot\psi^{n}(|x-z_{0i}|)\,dm(x)=\omega_{n-1}I_{i}\,$, (3.11)
where $I_{i}=\int\limits_{r^{i}_{1}}^{r^{i}_{2}}\frac{dt}{tq_{z_{0i}}^{\frac{1}{n-1}}(t)}.$ Note that, the function $\eta_{1}(t)=\psi(t)/I_{i},$ $t\in(r^{i}_{1},r^{i}_{2}),$ satisfies (1.4), because $\int\limits_{r^{i}_{1}}^{r^{i}_{2}}\eta_{1}(t)\,dt=1.$ Now, it follows from (3.11) that
$M(f_{m}(\Gamma(S(z_{0i},r^{i}_{1}),S(z_{0i},r^{i}_{2})),A(z_{0i},r^{i}_{1},r^{i}_{2})))\leqslant\frac{\omega_{n-1}}{I_{i}^{n-1}}<\infty\,$, (3.12)
where $I_{i}=I_{i}(z_{0i},r^{i}_{1},r^{i}_{2})=\int\limits_{r^{i}_{1}}^{r^{i}_{2}}\frac{dr}{rq_{z_{0i}}^{\frac{1}{n-1}}(r)}$ and $I_{i}<\infty$ because $q_{z_{0i}}(r)<\infty$ for $r\in E_{1}$ and some set of positive linear Lebesgue measure by the assumption of the lemma.
@@
$M(\Gamma(f_{m}(A_{m}),E_{m},\overline{{\mathbb{R}}^{n}}))\leqslant M(\Gamma_{m})$
$=M(f_{m}(\Delta_{m}))\leqslant\sum\limits_{i=1}^{n}M(f_{m}(\Gamma(S(z_{0i},r^{i}_{1}),S(z_{0i},r^{i}_{2}),A(z_{0i},r^{i}_{1},r^{i}_{2}))))$ (3.13)
@@
Since $P$ in (3.3) may be done arbitrary big, the relations (3.3) and (3.13) contradict each other. This completes the proof. $\Box$
@@ 4 Proof of the main results
By the compactness of $\overline{{\mathbb{R}}^{n}},$ we may consider that $x_{m},y_{m}\rightarrow x_{0}$ as $m\rightarrow\infty.$ There are two cases: 1) $x_{0}\in D_{0},$ 2) $x_{0}\in\partial D_{0}.$ In the first case, when $x_{0}$ is the inner point of $D_{0},$ all mappings $f_{m}$ except of finite numbers are defined in some ball $B(x_{0},\varepsilon_{0})\subset D_{0}\cap D_{m},$ $m\geqslant M_{0}.$ Now, $f_{m}$ is equicontinuous at $x_{0}$ (see [RS1, Theorems 6.1, 6.5] for $p=n$ and [Sev2, Theorems 4.1, 4.3] for n1<p<n.n-1<p<n. In this case, the right side in (4.2) tends to zero as $m\rightarrow\infty$ which contradicts with (4.1), and the proof finishes.
@@
It remains to show that $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots$ and some $\delta_{1}>0$ under conditions $p=n$ and 5), 6) of the theorem. We cover the continuum $A$ by the balls $B(x,r)$ with $x\in A$ and $0<r<\frac{1}{2}{\rm dist}\,(A,\partial D_{0}).$ Since $A$ is a compactum, we may choose a finite subcover $\bigcup\limits_{i=1}^{N}B(x_{i},r),$ $x_{i}\in A,$ of $A.$ Observe that, $h(\overline{B(x_{i},r)},\partial D_{m})\geqslant r_{*}>0$ for some $r_{*}>0,$ all $1\leqslant i\leqslant N$ and all $m\in{\mathbb{N}}$ except finite numbers. Otherwise, there are sequences $x_{m}\in\overline{B(x_{i},r)}$ and $y_{m}\in\partial D_{m}$ with $h(x_{m},y_{m})\rightarrow 0$ as $m\rightarrow\infty.$ We may consider that $x_{m},y_{m}\rightarrow x_{0}$ as $m\rightarrow\infty.$ Obviously, $x_{0}\in\overline{B(x_{i},r)}$ by the closeness of $\overline{B(x_{i},r)}.$ Now, there is $\varepsilon_{0}>0$ such that $\overline{B(x_{0},\varepsilon_{0})}\subset D_{0}$ and $\overline{B(x_{0},\varepsilon_{0})}$ belongs to all $D_{m}$ except of finite numbers. However, the latter contradicts the condition $y_{m}\rightarrow x_{0}$ as $m\rightarrow\infty$ while $y_{m}\in\partial D_{m}.$ Indeed, in this case $\overline{B(x_{0},\varepsilon_{0})}\cap\partial D_{m}\neq\varnothing$ for infinitely many $m\in{\mathbb{N}}.$
Finally, observe that $f_{m}\in\mathfrak{I}_{Q,\delta,N}(\mathfrak{D}),$ where $\mathfrak{D}=\{D_{m}\}_{m=1}^{\infty},$ $A_{f_{m}}=A.$ By Lemma 3, $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots$ and some $\delta_{1}>0,$ as required. $\Box$

### v2

@@ 3 On image of the continuum under open discrete and closed mappings
Indeed, let $\gamma:[0,1]\rightarrow\overline{{\mathbb{R}}^{n}}$ be a path in $\Gamma(f_{m}(A),E_{m},\overline{{\mathbb{R}}^{n}}),$ i.e., $\gamma(0)\in f_{m}(A),$ $\gamma(1)\in E_{m}$ and $\gamma(t)\in\overline{{\mathbb{R}}^{n}}$ for $t\in(0,1).$ Let $t_{m}=\sup\limits_{\gamma(t)\in f_{m}(D_{m})}t$ and let $\alpha_{m}(t)=\gamma|_{[0,t_{m})}.$ Let $\Gamma_{m}$ consists of all such paths $\alpha_{m},$ now $\Gamma(f_{m}(A),E_{m},\overline{{\mathbb{R}}^{n}})>\Gamma_{m}$ and by the minorization principle of the modulus (see [Fu, Theorem 1])
$M(\Gamma_{m})\geqslant M(\Gamma(f_{m}(A),E_{m},\overline{{\mathbb{R}}^{n}}))\,.$ (3.6)
Let $\Delta_{m}$ be a family of all total $f_{m}$ liftings of $\Gamma_{m}$ starting at $A$ (it exists by Proposition 3). Now
@@
We now show that
$\Delta_{m}>\bigcup\limits_{i=1}^{N}\Gamma(S(x_{i},\varepsilon_{i}),S(x_{i},2\varepsilon_{i}),A(x_{i},\varepsilon_{i},2\varepsilon_{i}))\,.$ (3.8)
@@
Indeed, let $\gamma\in\Delta_{m},$ in other words, $\gamma:[0,t_{m})\rightarrow D_{m},$ $\gamma(0)\in A$ and $\gamma(t)\rightarrow\partial D_{m}$ as $t\rightarrow t_{m}-0.$ Now, there is $1\leqslant i\leqslant N$ such that $|\gamma|\cap B(x_{i},\varepsilon_{i})\neq\varnothing\neq|\gamma|\cap(D_{m}\setminus B(x_{i},\varepsilon_{i})).$ Therefore, by Proposition 3 there is $0<t^{\,*}_{1}<t_{m}$ such that $\gamma(t^{\,*}_{1})\in S(x_{i},\varepsilon_{i}).$ We may assume that $\gamma(t)\not\in B(x_{i},\varepsilon_{i})$ for $t>t^{\,*}_{1}.$ Put $\gamma_{1}:=\gamma|_{[t^{\,*}_{1},t_{m}]}.$ Similarly, $|\gamma_{1}|\cap B(x_{i},2\varepsilon_{i})\neq\varnothing\neq|\gamma_{1}|\cap(D_{m}\setminus B(x_{i},2\varepsilon_{i})).$ By Proposition 3 there is $t^{\,*}_{1}<t^{\,*}_{2}<t_{m}$ with $\gamma(t^{\,*}_{2})\in S(x_{i},2\varepsilon_{i}).$ We may assume that $\gamma(t)\in B(x_{i},\varepsilon_{i})$ for $t<t^{\,*}_{2}.$ Put $\gamma_{2}:=\gamma|_{[t^{\,*}_{1},t^{\,*}_{2}]}.$ Then, the path $\gamma_{2}$ is a subpath of $\gamma,$ which belongs to the family $\Gamma(S(x_{i},\varepsilon_{i}),S(x_{i},2\varepsilon_{i}),A(x_{i},\varepsilon_{i},2\varepsilon_{i})).$ Thus, the relation (3.8) is established.
@@
$\psi_{i}(t)=\left\{\begin{array}[]{rr}1/[tq_{x_{i}}^{\,\prime\frac{1}{n-1}}(t)]\ ,&\ t\in(\varepsilon_{i},2\varepsilon_{i})\ ,\\ 0\ ,&\ t\notin(\varepsilon_{i},2\varepsilon_{i})\ ,\end{array}\right$.
where $q^{\,\prime}_{x_{i}}(r):=\frac{1}{\omega_{n-1}r^{n-1}}\int\limits_{|x-x_{i}|=r}Q^{\,\prime}(x)\,dS,$ $Q^{\,\prime}(x)=\max\{Q(x),1\}.$ Now,
$\int\limits_{A(x_{i},\varepsilon_{i},2\varepsilon_{i})}Q^{\,\prime}(x)\cdot\psi^{n}(|x-x_{i}|)\,dm(x)=\omega_{n-1}I_{i}\,$, (3.9)
where $I_{i}=\int\limits_{\varepsilon_{i}}^{2\varepsilon_{i}}\frac{dt}{tq_{x_{i}}^{{}^{\,\prime}\frac{1}{n-1}}(t)}.$ Note that, the function $\eta_{i}(t)=\psi(t)/I_{i},$ $t\in(\varepsilon_{i},2\varepsilon_{i}),$ satisfies (1.4), because $\int\limits_{\varepsilon_{i}}^{2\varepsilon_{i}}\eta_{i}(t)\,dt=1.$ Now, it follows from (3.9) that
$M(f_{m}(\Gamma(S(x_{i},\varepsilon_{i}),S(x_{i},2\varepsilon_{i}),A(x_{i},\varepsilon_{i},2\varepsilon_{i})))\leqslant\frac{\omega_{n-1}}{I_{i}^{n-1}}<\infty\,$, (3.10)
where $I_{i}=I_{i}(x_{i},\varepsilon_{i},2\varepsilon_{i})=\int\limits_{\varepsilon_{i}}^{2\varepsilon_{i}}\frac{dr}{rq_{x_{i}}^{\,\prime\frac{1}{n-1}}(r)}$ and $I_{i}>0$ by the construction.
@@
$M(\Gamma(f_{m}(A),E_{m},\overline{{\mathbb{R}}^{n}}))\leqslant M(\Gamma_{m})$
$=M(f_{m}(\Delta_{m}))\leqslant\sum\limits_{i=1}^{N}M(f_{m}(\Gamma(S(x_{i},\varepsilon_{i}),S(x_{i},2\varepsilon_{i}),A(x_{i},\varepsilon_{i},2\varepsilon_{i}))))$ (3.11)
@@
Since $P$ in (3.5) may be done arbitrary big, the relations (3.5) and (3.11) contradict each other. This completes the proof for the case when the relation (1.6) holds. Observe that, the case when $Q$ has a finite mean oscillation in $A$ is a particular case of the relation (1.6) (see the last part of the proof of Theorem 5.5 in [Sev2]). The proof is wholly complete. $\Box$
Corollary 3.1.   Assume that, under conditions of Lemma 3, $A_{0}$ is some another compactum in $D_{0}.$ Then there is $\delta_{2}>0$ and $M_{0}\in{\mathbb{N}}$ such that $A_{0}\subset D_{m}$ for all $m\geqslant M_{0}$ and $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0}.$
Proof.   Since $D_{m}$ converges to $D_{0}$ as its kernel, there is $M_{0}\in{\mathbb{N}}$ such that $A_{0}\subset D_{m}$ for all $m\geqslant M_{0}.$ It remains to show that $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0}$ and some $\delta_{2}>0.$ Assume the contrary, i.e., $h(f_{m}(A_{0}),\partial f_{m}(D_{m})\rightarrow 0$ as $m\rightarrow\infty.$ Now, $h(f_{m}(A_{0}),\partial f_{m}(D_{m})=h(x_{m},y_{m})\rightarrow 0$ as $m\rightarrow\infty$ for $x_{m}\in A_{0}$ and $y_{m}\in f_{m}(D_{m}).$ We may consider that $x_{m}\rightarrow x_{0}$ as $m\rightarrow\infty,$ $x_{0}\in A_{0}.$ Put $\varepsilon_{0}>0$ such that $\overline{B(x_{0},\varepsilon_{0})}\subset D_{0}.$ We join $x_{0}$ with $A$ by a path $\gamma$ in $D_{0}.$ Now, $B:=A\cup|\gamma|\cup\overline{B(x_{0},\varepsilon_{0})}$ is a continuum in $D_{0}$ such that $B\subset D_{m}$ for sufficiently large $m,$ $x_{m}\in B$ for sufficiently large $m$ and $h(f_{m}(B))\geqslant h(f_{m}(B))\delta$ for above $m.$ By Lemma 3, $h(f_{m}(B),\partial f_{m}(D_{m}))\geqslant\delta_{*}>0$ for $m\geqslant m_{3},$ $m_{3}\in{\mathbb{N}}.$ However, the latter contradicts with $h(f_{m}(A_{0}),\partial f_{m}(D_{m})=h(x_{m},y_{m})\rightarrow 0$ as $m\rightarrow\infty,$ because obviously $h(f_{m}(B),\partial f_{m}(D_{m}))\leqslant h(x_{m},y_{m})\rightarrow 0$ as $m\rightarrow\infty.$ The obtained contradiction completes the proof. $\Box$
@@ 4 Proof of the main results
By the compactness of $\overline{{\mathbb{R}}^{n}},$ we may consider that $x_{m},y_{m}\rightarrow x_{0}$ as $m\rightarrow\infty.$ There are two cases: 1) $x_{0}\in D_{0},$ 2) $x_{0}\in\partial D_{0}.$ In the first case, when $x_{0}$ is the inner point of $D_{0},$ all mappings $f_{m}$ except of finite numbers are defined in some ball $B(x_{0},\varepsilon_{0})\subset D_{0}\cap D_{m},$ $m\geqslant M_{0}.$ Now, $f_{m}$ is equicontinuous at $x_{0}$ (see [RS, Theorems 6.1, 6.5] for $p=n$ and [Sev2, Theorems 4.1, 4.3] for n1<p<n.n-1<p<n. In this case, the right side in (4.2) tends to zero as $m\rightarrow\infty$ which contradicts with (4.1), and the proof finishes.
@@
The existing of $\delta_{1}>0$ for which $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots$ follows by Lemma 3. The existing of $\delta_{2}>0$ for which $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for any $m=1,2,\ldots$ follows by Corollary 3.  $\Box$